# Multivariate Causal Effects: a Bayesian Causal Regression Factor Model

### v1

@@ 3. Bayesian Causal Factor Model
By the the properties of the multivariate normal distribution, the conditional distribution of the factor scores given the covariates and the unmeasured variable is given by
@@
where the location and scale parameters are defined as following
@@
Although the conditional distribution in Equation (15) cannot be directly estimated due to the unmeasured variable due to the unmeasured variable $U$, the following property 2 demonstrates that our proposed formulation for the distribution of treatment-specific factors ${\bf L}_{t}$ is sufficient to recover the relevant information about $U$.
@@
Assuming the unmeasured variable $U$ is normal distributed, the treatment-specific factors ${\bf L}_{t}$, for each $t$, are defined as in (3)-(5), and the assumptions of Eq.(11) hold, then each component of the mixture $r\geq 1$, has: (i) a finite expected value of the location parameter, that takes into account the expected value of the unmeasured variable $U$, and (ii) the scale parameter has the scale parameter of the conditional distribution (15) as a lower bound. Specifically,
@@
Therefore, the probability distribution of $\{L_{tj}\mid{\bf X}\}$ has same mean of $\{L_{tj}\mid{\bf X},U\}$ and an equal or larger variance. Substituting this result into Eq. (1), we conclude that the casual effects—the SATE—are unbiased under the assumptions defined in Section 2.
Scenario 1 in Figure 1 corresponds to the case where factor scores and the unmeasured variable are independent, i.e., $\sigma^{(lu)}=0$. In this case, the marginal and conditional variances are the same, indeed the equation (16) can be simplified in $1/\tau_{r}^{(tj)}=\sigma^{(l|u)2}_{r}$. Conversely, Scenarios 2 and 3 assume $\sigma^{(lu)}>0$. Assuming a positive value for $\sigma^{(lu)}$ does not determinate the direction of dependence, thus encompassing both cases.
@@
Alternative priors include the Bayesian cumulative shrinkage introduced by Legramanti et al. (2020) or a generalization of Bhattacharya and Dunson (2011)’s prior proposed by Schiavon et al. (2022).
@@ 4. Simulation Study
We evaluate the performance of the proposed model through an extensive simulation study. The primary objective is to assess its ability to estimate causal effects, with a particular focus on controlling the variability of SATE and adjusting the bias induced by the unmeasured variable $U$. We compare our causal factor model against state-of-the-art flexible models in causal inference—causal BART by Hill (2011) and Bayesian causal forest (BCF) introduced by Hahn et al. (2020)—as well as a standard factor model employing a standard normal prior for factor scores and adjusted for the causal inference setting. Specifically, the standard factor model is defined as in (2), but with priors ${\bf l}_{it}\sim\mathcal{N}(0,1)$ for each unit $i$ and treatment level $t\in\{0,1\}$. This setup allows us to compare and evaluate the impact of the proposed DDP prior on treatment-specific factor scores. While BART and BCF are estimated separately for each element of the outcome variable, both the standard factor model adjusted for causal inference and our proposed approach explicitly account for the multivariate nature of the outcome.
We consider four distinct simulation scenarios. The first three explore different relationships between the unmeasured variable $U$ and the confounders ${\bf X}$, corresponding to the graphical representations in Figure 1. The fourth scenario is designed to closely mimic the real dataset used in our application in Section 5. Details of simulation-generating processes are provided in Table 1.
@@
Specifically, in the first three scenarios, the treatment-specific factor scores $\{{\bf L}_{0},{\bf L}_{1}\}$ depend on both the confounders ${\bf X}$ and the unmeasured variable $U$, while the potential outcomes ${\bf Y}$ depend on $U$ only through the factors. In Scenario 1, illustrated on the left of Figure 1, ${\bf X}$ and $U$ are independent (i.e., $\sigma^{lu}=0$ in Property 2). Scenario 2 assumes $U$ depends on ${\bf X}$, and Scenario 3 explores the reverse relationship, where the confounders ${\bf X}$ depend on $U$. These dependencies are highlighted by the red dotted arrows in the second and third representations of Figure 1. To simulate that, as reported in Table 1, Scenario 2 has four independent confounders ${\bf X}$ and $U$ dependent on them, while Scenario 3 is defined with an independent normal distribution for $U$ and the first two confounders ${\bf X}$ as normal distributed with mean regression dependent of $U$. Scenario 4 uses real-data information for the 28 confounders and the treatment assignment. Factor loadings are estimates with our model applied to real data. Factor scores and outcomes were simulated according to our proposed model.
@@
We evaluate the performance and accuracy of the models according to the bias mean and square error (MSE) of SATEk  estimation for each $k=\{1,\dots,10\}$ in Scenarios 1–3, and $k=\{1,\dots,27\}$ in Scenario 4. The results are summarized in Figure 2. Our proposed model (CausalFA) outperforms the competitors in both bias and MSE, showing unbias estimations of the causal effects and smaller variability. Specifically, BART and BCF exhibit an overestimation of variability, leading to high MSE values, likely due to their independent estimation of each causal effect. The factor model with the standard normal prior (StandardFA) reveals bias estimation in the causal effects, highlighting a critical weakness of the standard prior for the factor scores and the benefit of the DDP prior.
Our proposed model is the only one among the four implemented models to provide an unbiased estimation of the causal effects for all the elements in the multivariate outcomes across the simulated scenarios, demonstrating its ability to adjust the estimation for the unmeasured variable through the latent factors and their prior.
@@ 5. Environmental Application
Following Krasovich Southworth et al. (2025), who report strong temporal correlations in the data, we focus on a specific time period—specifically, the summer months—when wildfires are more frequent. Moreover, in their analysis they highlight the difficulties in defining the level of pm2.5 induced by wildfire smoke and of other origins due to different measurement technologies. To account for this, our treatment variable is defined as the binary presence or absence of wildfire smoke.
@@
Each monitor records daily information on the level of each chemical species, which is collected every three days. The IMPROVE and CSN monitoring data were preprocessed to replace data flagged as unacceptable quality with interpolated values or, when appropriate, a value equal to 1/2 of the minimum detection limit (less than 3% of values were interpolated). To adjust for methodological differences between the CSN and IMPROVE monitors, we include the variable monitor_type as a confounder.
Wildfire smoke pm2.5 concentration are derived from the daily prediction of Childs et al. (2022), available from the GitHub repository echolab-stanford/daily-10km-smokePM. We define the treatment as $T=1$ for “smoke days” with non-zero predictions and control $T=0$ for “non-smoke days” with zero predicted wildfire smoke pm2.5. The prediction resolution is in the grid $10Km$ (i.e., each unit aggregates the information over $10Km\times 10Km$ area), therefore we match the location of each air pollution monitor with the closest point of the grid.
As confounders, we collect weather information (from the Harvard Dataverse (Audirac, 2024)) for U.S. ZIP Code Tabulation Areas (ZCTA) at daily resolution and census data (via the tidycensus R package (Walker et al., 2021)). Weather variables include daily values for near-surface air temperature, maximum near-surface air temperature, precipitation, minimum and maximum near-surface relative humidity, surface downwelling solar radiation, and wind speed and direction at $10m$. Census variables include population, percentage of males, percentage of different ethnicities (White, Black and African American, Asian, and other), median household income, total number of housing units, and poverty status. Both the weather information and census data are matched with the monitor locations.
@@
The treatment variable is defined as wildfire smoke exposure $T=0$ for non-smoke days and $T=1$ for smoke days. The multivariate outcome is the concentration of $27$ chemical species, measured in $\mu g/m^{3}$ and transformed on the natural logarithmic scale. Confounders include weather information, census data, type of monitor indication, month, latitude and longitude of monitor location, and type of monitor location (rural, city or intermediate).
Initially we have $5,754$ monitor-day units assigned in the control group—that is, no wildfire smoke day in that monitor location—and $1,713$ in the treated group—that is, exposed to wildfire smoke. However, we use matching before running our model to make our analyses as robust as possible with respect to the potential measured confounding bias, using confounders to measure the similarity of the units in the two groups. This is common in observational studies (Rosenbaum and Rubin, 1983) and in research on air pollution effects on health (see, e.g., Lee et al., 2021; Zorzetto et al., 2024).
@@ Results
Similarly to the simulation study, we estimate the causal effect of wildfire smoke on $27$ chemical species using four models: our proposed causal factor model, the factor model with a standard Gaussian distribution for the factor score, BART, and BCF. The results are compared in Figure 3.
All four models consistently identify a positive causal effect of wildfire smoke on zinc, bromine, and the two organic elements—elemental and organic carbon—indicating that wildfire smoke increases the levels of these chemicals. This finding aligns with previous studies: Young and Jan (1977) and Odigie and Flegal (2014) observed increased zinc concentrations following wildfire smoke in California, Liu et al. (2014) reported elevated carbon levels and their negative impact on ecosystems, and Li et al. (2023) found higher bromine concentrations during wildfires in urban areas. However, the models do not fully agree on the detection of significant negative causal effects, i.e. a reduction in chemical concentrations due to wildfire smoke, particularly for strontium, sodium, nickel, titanium, silicon and aluminum. This discrepancy arises from the wide credible intervals produced by BART, which often include zero, and, in some cases, by BCF as well, limiting detection of significant effects. As shown in the simulation study, these two models tend to overestimate data variability, particularly in highly correlated outcomes. Similarly for both positive effects, only our proposed model has credible intervals that do not include the zero for cooper, lead, sulfate, and phosphorum. This is consistent with previous studies, as Young and Jan (1977) and Odigie and Flegal (2014) observed notable increases in these metals, while Spencer and Hauer (1991) reported elevated phosphorus concentrations following wildfires.
@@
The varying behaviors identified by the four models can be better understood by examining the factor loading matrix, based on correlations among the 27 chemical species. Our Bayesian causal factor model framework enables us to analyze these relationships and assess how wildfire smoke alters them by estimating treatment-specific factors—i.e., distinct factors for the presence and absence of wildfire smoke. After applying a varimax transformation (Kaiser, 1960), we identify $3$ factors for each treatment level, explaining a total of $85\%$ and $88\%$ of the variance, respectively for wildfire smoke and non-smoke days (Figure 4). In both treatment levels, the first factor captures strong correlations among transition metals, metalloids, and other metals. Notably, titanium, silicon, and aluminum—chemicals whose concentrations are found to be strongly correlated by Factor 1—are detected to have significant wildfire smoke effects by only the factor models, highlighting their ability to include correlation across outcomes and narrow uncertainty and more robust conclusions.
Factor 2 captures strong correlations among nonmetals, while Factor 3 reflects associations among organic elements. In both cases, wildfire smoke is estimated to increase their levels. This can be attributed to wildfires burning natural materials such as forests, releasing elemental and organic carbon into the air Liu et al. (2014). Likewise, nonmetals, commonly found in soil, exhibit strong positive correlations with wildfire events (Spencer and Hauer, 1991).
Furthermore, the presence of wildfire smoke appears to enhance correlations among alkaline-earth metals, transition metals, and metalloids within Factor 1, particularly among chemicals that share the same negative causal effect. Similarly, in Factor 2, wildfire smoke strengthens correlations among nonmetals, while in Factor 3, it disrupts the correlation between organic chemicals and other elements, reinforcing their internal associations. These structural shifts underscore the utility of our model in uncovering the latent structure of treatment.
@@ 6. Discussion
This work introduces two key innovations. First, from a factor analysis perspective, we propose an infinite mixture distribution for the factor scores, addressing a critical limitation in the existing literature, which has historically focused on factor loadings while largely overlooking the role of latent scores. Second, within the causal inference framework, we develop a multivariate outcome setup that enables the sharing of latent structure across outcome components, allowing for the recovery of unmeasured variables and improving the estimation of causal effects.
@@
By combining these two contributions, the proposed Bayesian causal regression factor model demonstrates superior performance in causal effect estimation compared to state-of-the-art methods in both the factor analysis and causal inference domains, also correcting the bias in the causal effect estimates. Specifically, it outperforms standard factor models that rely on Gaussian priors for factor scores, which fail to capture heterogeneity and impute missing data. At the same time, it outperforms established methods in causal inference such as BART and BCF, which, despite their flexibility, lack the ability to share latent structure across correlated outcome components, leading to uncertainty and exhibiting weaknesses in handling variability.
@@
While the proposed approach can be broadly applied to various real-world datasets for studying the effects of treatments or exposures on outcomes composed of correlated elements, our application in this paper focuses on a critical environmental issue. Specifically, we aim to disentangle the causal effect of wildfire smoke on the chemical composition of pm2.5, driven by awareness of the significant role that fine particulate exposure plays in human health. Our findings not only identify the chemicals that are positively or negatively affected by wildfire smoke but also reveal significant effects that other models fail to detect, due to its ability to borrow strength across outcomes. Moreover, our method effectively identifies the latent factors that drive the behavior of chemical components across different types of metals and non-metals, capturing treatment-specific patterns across classes of metals, nonmetals, and organics. This capability enriches our understanding of how wildfire smoke alters the composition of air quality in complex and systematic ways.
@@
Our work aims to serve as a foundation for a growing body of research that highlights the importance of factor scores—supported by recent studies such as Zorzetto et al. (2024) and Bortolato and Canale (2024)—and explores the potential of factor models to quantify treatment effects in causal inference, particularly in settings where outcomes are high-dimensional and strongly correlated.

### v2

@@ 3. Bayesian Causal Factor Model
From the properties of the multivariate normal distribution, the conditional distribution of the factor scores given the covariates and the unmeasured variable is
@@
where the location and scale parameters are
@@
Although the conditional distribution in Equation (15) cannot be directly estimated as $U$ is not measured, the following property 2 demonstrates that our proposed formulation for the distribution of treatment-specific factors ${\bf L}_{t}$ is sufficient to recover the information about $U$ on average.
@@
Assuming the unmeasured variable $U$ is normally distributed, the treatment-specific factors ${\bf L}_{t}$, for each $t$, are defined as in (3)-(5), and the assumptions of Eq.(11) hold, then each component of the mixture $r\geq 1$, has: (i) a finite expected value of the location parameter, that depend on the expected value of the unmeasured variable $U$, and (ii) the scale parameter has the scale parameter of the conditional distribution (15) as a lower bound. Specifically,
@@
Therefore, the probability distribution of $\{L_{tj}\mid{\bf X}\}$ has same mean as $\{L_{tj}\mid{\bf X},U\}$ and an equal or larger variance. Substituting this result into Eq. (1), we conclude that the casual effects—the SATE—are unbiased under the assumptions defined in Section Causal Estimands and Identifying Assumptions.
Scenario 1 in Figure 1 corresponds to the case where factor scores and the unmeasured variable are independent, i.e., $\sigma^{(lu)}=0$. In this case, the marginal and conditional variances are the same, indeed the equation (16) can be simplified in $1/\tau_{r}^{(tj)}=\sigma^{(l|u)2}_{r}$. Conversely, Scenarios 2 and 3 assume $\sigma^{(lu)}>0$.
@@
We follow the default hyperparameter settings proposed by Bhattacharya and Dunson (2011).
@@ 4. Simulation Study
We evaluated the performance of the proposed model through an extensive simulation study. The primary objective is to assess its ability to estimate causal effects, with a particular focus on controlling the variability of SATE and adjusting the bias induced by the unmeasured variable $U$. We compare our causal factor model to state-of-the-art flexible models in causal inference—causal BART by Hill (2011) and Bayesian causal forest (BCF) by Hahn et al. (2020)—as well as a standard factor model employing a standard normal prior for factor scores and adapted the causal inference setting. Specifically, the standard factor model is defined as in (2), but with priors ${\bf l}_{it}\sim\mathcal{N}(0,1)$ for each unit $i$ and treatment level $t\in\{0,1\}$. This setup allows us to compare and evaluate the impact of the proposed DDP prior on treatment-specific factor scores. While BART and BCF are estimated separately for each element of the outcome variable, both the standard factor model adapted to causal inference and our proposed approach explicitly account for the multivariate nature of the outcome.
We consider four simulation scenarios. The first three explore different relationships between the unmeasured variable $U$ and the confounders ${\bf X}$, corresponding to the graphical representations in Figure 1. The fourth scenario is designed to closely mimic the real dataset used in our application in Section 5. Environmental Application. Details of simulation-generating processes are provided in Table 1.
@@
Specifically, in the first three scenarios, the treatment-specific factor scores $\{{\bf L}_{0},{\bf L}_{1}\}$ depend on both the confounders ${\bf X}$ and the unmeasured variable $U$, while the potential outcomes ${\bf Y}$ depend on $U$ only through the factors. In Scenario 1, shown on the left of Figure 1, ${\bf X}$ and $U$ are independent (i.e., $\sigma^{lu}=0$ in Property 2). Scenario 2 assumes that $U$ depends on ${\bf X}$, and Scenario 3 explores the reverse relationship, where the confounders ${\bf X}$ depend on $U$. These dependencies are highlighted by the red dotted arrows in the second and third representations of Figure 1. Accordingly to Scenario 2, we consider four independent confounders ${\bf X}$ and model $U$ as dependent on ${\bf X}$. In contrast, Scenario 3 is modeled with an independent normal distribution for $U$; also the first two confounders in ${\bf X}$ are normally distributed with mean given by a linear function of $U$. Scenario 4 uses real-data information for the 28 confounders and the treatment assignment. Factor loadings are estimated with our model applied to real data. Factor scores and outcomes were simulated according to our proposed model.
@@
We evaluate the performance of the models according to the bias mean and square error (MSE) of SATEk  estimation for each $k=\{1,\dots,10\}$ in Scenarios 1–3, and $k=\{1,\dots,27\}$ in Scenario 4. The results are summarized in Figure 2. Our proposed model (CausalFA) outperforms the competitors in both bias and MSE, showing unbiased estimations of causal effects and a smaller variability. Specifically, BART and BCF exhibit higher values for the MSE, with particular ample boxplots for the BART. The factor model with standard normal prior (StandardFA) estimates causal effects with bias, highlighting a critical weakness of standard prior for factor scores and the benefit of the DDP prior.
Our proposed model is the only one among the four we considered to provide an unbiased estimation of the causal effects for all the elements in the multivariate outcomes across the simulated scenarios, demonstrating its ability to adjust the estimation for the unmeasured variable through the latent factors and their prior.
@@ 5. Environmental Application
Following Krasovich Southworth et al. (2025), who report strong temporal correlations in the data, we focus on a specific time period—specifically, the summer months—when wildfires are more frequent. Moreover, instead of using the prediction of the level of pm2.5 induced by wildfire smoke, our treatment variable is defined as the binary presence or absence of wildfire smoke.
@@
Each monitor records daily information on the level of each chemical species. The dataset includes data for one out of every three days. The IMPROVE and CSN monitoring data were preprocessed to replace data flagged as unacceptable quality with interpolated values or, when appropriate, a value equal to 1/2 of the minimum detection limit (less than 3% of values were interpolated). To adjust for methodological differences between the CSN and IMPROVE monitors, we include the variable monitor_type as a confounder.
Wildfire smoke pm2.5 concentration are derived from the daily prediction of Childs et al. (2022), available from the GitHub repository echolab-stanford/daily-10km-smokePM. We define the treatment variable as $T=1$ for “smoke days” with non-zero predictions and $T=0$ for “non-smoke days” with zero predicted wildfire smoke pm2.5. The prediction resolution in the grid is $10Km$ (i.e., each unit aggregates the information over $10Km\times 10Km$ area), therefore we match the location of each air pollution monitor with the closest point of the grid.
As confounders, we collect weather information (from the Harvard Dataverse (Audirac, 2024)) for U.S. ZIP Code Tabulation Areas (ZCTA) at daily resolution and census data (via the tidycensus R package (Walker et al., 2021)). Weather variables include daily values for near-surface air temperature, maximum near-surface air temperature, precipitation, minimum and maximum near-surface relative humidity, surface downwelling solar radiation, and wind speed and direction at $10m$. Census variables include population, percentage of males, percentage of different ethnicities (White, Black and African American, Asian, and other), median household income, total number of housing units, and poverty status. Both weather information and census data are matched with the monitor locations.
@@
The multivariate outcome is the concentration of $27$ chemical species, measured in $\mu g/m^{3}$ and transformed on the natural logarithmic scale. Confounders include weather information, census data, type of monitor indication, month, latitude and longitude of monitor location, and type of monitor location (rural, city or intermediate).
Initially we have $5,754$ monitor day units assigned in the control group —-that is, no wildfire smoke day at that monitor location —-and $1,713$ in the treated group —that is, exposed to wildfire smoke. However, we use matching before running our model to make our analyses as robust as possible with respect to the potential measured confounding bias, using confounders to measure the similarity of the units in the two groups. This is common in observational studies (Rosenbaum and Rubin, 1983) and in research on air pollution effects on health (see, e.g., Lee et al., 2021; Zorzetto et al., 2024).
@@ Results
Similarly to the simulation study, we estimate the causal effect of wildfire smoke on $27$ chemical species using four models: our proposed causal factor model, the factor model with a standard Gaussian distribution for the factor scores, BART, and BCF. The results are compared in Figure 3.
All four models consistently identify a positive causal effect of wildfire smoke on zinc, bromine, and the two organic elements—elemental and organic carbon—indicating that wildfire smoke increases the levels of these chemicals. This finding aligns with previous studies: Young and Jan (1977) and Odigie and Flegal (2014) observed increased zinc concentrations following wildfire smoke in California, Liu et al. (2014) reported elevated carbon levels and their negative impact on ecosystems, and Li et al. (2023) found higher bromine concentrations during wildfires in urban areas. However, the models do not fully agree on the detection of significant negative causal effects, i.e. a reduction in chemical concentrations due to wildfire smoke, particularly for strontium, sodium, nickel, titanium, silicon and aluminum. This discrepancy arises from the wide credible intervals produced by BART, which often include zero, and, in some cases, by BCF as well, limiting detection of significant effects. As shown in the simulation study, these two models tend to estimate wide variability, particularly in highly correlated outcomes. Similarly, for positive effects, only our proposed model has credible intervals that do not include zeros for cooper, lead, sulfate, and phosphorum. This is consistent with previous studies, as Young and Jan (1977) and Odigie and Flegal (2014) observed notable increases in these metals, while Spencer and Hauer (1991) reported elevated phosphorus concentrations following wildfires.
@@
The different distributions of the causal effects identified by the four models can be better understood by examining the factor loading matrix, reflecting the correlations among the 27 chemical species. Our Bayesian causal factor model framework enables us to analyze these relationships and assess how wildfire smoke alters them by estimating treatment-specific factors—i.e., distinct factors for the presence and absence of wildfire smoke. After applying a varimax transformation (Kaiser, 1960), we identify $3$ factors for each treatment level, explaining a total of $85\%$ and $88\%$ of the variance, respectively for wildfire smoke and non-smoke days (Figure 4). In both treatment levels, the first factor captures strong correlations among transition metals, metalloids, and other metals. Notably, titanium, silicon, and aluminum, chemicals whose concentrations drive the Factor 1, are detected to have significant wildfire smoke effects by only the factor models (CausalFA and StandardFA), highlighting their ability to include correlation across the outcome components and narrow uncertainty and more robust conclusions.
Factor 2 captures strong correlations among non-metals, while Factor 3 reflects associations among organic elements. In both cases, wildfire smoke is estimated to increase factor levels. This may be plausibly related to wildfires burning natural materials such as forests, releasing elemental and organic carbon into the air Liu et al. (2014). Likewise, non-metals, commonly found in soil, exhibit strong positive correlations with wildfire events (Spencer and Hauer, 1991).
Furthermore, the presence of wildfire smoke appears to enhance correlations among alkaline-earth metals, transition metals, and metalloids within Factor 1, particularly among chemicals that share the same negative causal effect. Similarly, in Factor 2, wildfire smoke strengthens correlations among non-metals, while in Factor 3, it disrupts the correlation between organic chemicals and other elements, reinforcing their internal associations. These structural shifts underscore the utility of our model in uncovering the latent structure of treatment.
@@ 6. Discussion
This work introduces two key innovations. First, from a factor analysis perspective, we propose an infinite mixture distribution for the factor scores, addressing a critical limitation in the existing literature that has historically focused on factor loadings while largely overlooking the role of latent factor scores. Second, within the causal inference framework, we develop a multivariate outcome setup that enables the sharing of latent structure across outcome components, allowing adjustment for unmeasured variables potentially affecting components of the latent structure. We showed how this can improve the estimation of vector-valued causal effects.
@@
By combining these two contributions, the proposed Bayesian causal regression factor model demonstrates superior performance in causal effect estimation compared to state-of-the-art methods in both the factor analysis and causal inference domains, also correcting the bias in the causal effect estimates. Specifically, it outperforms standard factor models that rely on Gaussian priors for factor scores, which fail to characterize heterogeneity in factor scores and consequently impute the missing outcome. At the same time, it outperforms established methods in causal inference such as BART and BCF, which, despite their flexibility, lack the ability to share latent structure across correlated outcome components. Moreover, BART and BCF exhibit higher MSE values, likely due to their independent estimation of each causal effect. However, future extensions could enhance our causal factor model by leveraging the flexibility of BART priors Chipman et al. (2010). Additionally, alternative priors can be considered for both the covariate regression—such as mixtures or spike-and-slab priors George and McCulloch (1993)—and the factor loadings, including the Bayesian cumulative shrinkage prior by Legramanti et al. (2020) and the generalized sparse prior of Schiavon et al. (2022).
@@
While the proposed approach can be broadly applied to various real-world datasets for studying the effects of treatments or exposures on outcomes composed of correlated elements, our application in this paper focuses on a critical environmental issue. Specifically, we aim to estimate the causal effect of wildfire smoke on the chemical composition of pm2.5, motivated by the significant role that fine particulate exposure plays in human health. Our findings not only identify the chemicals that are positively or negatively affected by wildfire smoke but also reveal significant effects that other models fail to detect, due to its ability to borrow strength across outcomes. Moreover, our method effectively identifies the latent factors that drive the behavior of chemical components across different types of metals and non-metals, capturing treatment-specific patterns across classes of metals, nonmetals, and organics. This capability enriches our understanding of how wildfire smoke alters the composition of air quality in complex and systematic ways. Further research could extend this work by modeling multivariate outcomes under exposure to wildfire smoke over time, with methodologies similar to Samartsidis et al. (2020, 2024) or by accounting for the positive continuous nature of the prediction of pm2.5 induced by wildfire smoke, as discussed in Krasovich Southworth et al. (2025), though in a non-causal framework.
@@
We hope that this study highlights the importance of factor scores—not only in causal inference, as supported by recent studies such as Zorzetto et al. (2024) and Bortolato and Canale (2024)—and indicates the potential of factor models to quantify treatment effects in causal inference, particularly in settings where outcomes are high-dimensional and strongly correlated.